# Multimodal Inconsistency Reasoning (MMIR): A New Benchmark for Multimodal Reasoning Models

### v2

@@ Abstract
Existing Multimodal Large Language Models (MLLMs) are predominantly trained and tested on consistent visual-textual inputs, leaving open the question of whether they can handle inconsistencies in real-world, layout-rich content. To bridge this gap, we propose the Multimodal Inconsistency Reasoning (MMIR) benchmark to assess MLLMs’ ability to detect and reason about semantic mismatches in artifacts such as webpages, presentation slides, and posters. MMIR comprises 534 challenging samples, each containing synthetically injected errors across five reasoning-heavy categories: Factual Contradiction, Identity Misattribution, Contextual Mismatch, Quantitative Discrepancy, and Temporal/Spatial Incoherence. We evaluate six state-of-the-art MLLMs, showing that models with dedicated multimodal reasoning capabilities, such as o1, substantially outperform their counterparts while open-source models remain particularly vulnerable to inconsistency errors. Detailed error analyses further show that models excel in detecting pairwise inconsistencies but struggle with inconsistencies confined to single elements in complex layouts. Probing experiments reveal that single-modality prompting, including Chain-of-Thought (CoT) and Set-of-Mark (SoM) methods, yields marginal gains, revealing a key bottleneck in cross-modal reasoning. Our findings highlight the need for advanced multimodal reasoning and point to future research on multimodal inconsistency.
@@ 1 Introduction
Recent advances in Large Language Models (LLMs) have demonstrated impressive reasoning abilities across a variety of tasks OpenAI (2024b); Guo et al. (2025); Kojima et al. (2022); Wei et al. (2022). Building on pre-trained LLMs, Multimodal Large Language Models (MLLMs) are fast evolving. However, they usually face greater challenges as they need to reason across different modalities, especially when inconsistencies (i.e., mismatched or contradictory contents) exist. We find that, being primarily trained and evaluated on consistent visual-textual inputs, existing MLLMs are largely untested in scenarios where the input contains misaligned or contradictory information—a situation that is common in real-world scenarios. For example, in Figure 1, a user presents a web page containing conflicting visual and textual elements, asking the model to identify errors.
@@
To comprehensively evaluate the ability of MLLMs in reasoning over multimodal inconsistency, we introduce the Multimodal Inconsistency Reasoning Benchmark (MMIR). MMIR is the first framework dedicated to evaluating how effectively MLLMs can reason about and identify semantic mismatches within complex, layout-rich content with interleaved image and text components. Our benchmark is built on a diverse collection of real-world artifacts (e.g. websites, slides, posters) which have been augmented with synthetic inconsistencies—realistic inconsistency errors injected into their original structures. These inconsistency errors span a range of reasoning-heavy categories: Factual Contradiction, Identity Misattribution, Contextual Mismatch, Quantitative Discrepancy, and Temporal/Spatial Incoherence—posing a next-level reasoning challenge for models. For example, resolving a Identity Misattribution involves verifying entity alignment across modalities, while Quantitative Discrepancy requires cross-referencing chart data with textual claims. By challenging models to detect such inconsistencies, MMIR forces them to perform intricate reasoning that goes well beyond simple pattern recognition. This benchmark not only exposes the limitations of current MLLMs in handling real-world challenges of reasoning over multimodal content with inconsistency, but also provides a platform for developing more robust multimodal reasoning systems.
@@
In our experiments, we evaluated the advanced multimodal reasoning model o1 OpenAI (2024b) and five other state-of-the-art MLLMs: GPT-4o OpenAI (2024a), Qwen2.5-VL Team (2025), LLaVA-NeXT Liu et al. (2024b), InternVL2.5 Chen et al. (2024) and Phi-3.5-Vision Abdin et al. (2024) using MMIR’s 534 test samples. The results overall underscore that current MLLM models struggle with multimodal inconsistency reasoning. Specifically, there is a stark contrast between proprietary and open-source models. The open-source models evaluated only reach less than 25% accuracy. o1 with strong reasoning capability achieves the overall best performance with over 50% accuracy.
@@
To further understand the benchmarking results, we conduct analysis based on the inconsistency category, modality, and layout complexity of the artifact. We find the proprietary models excel in identifying factual contradiction and identity misattribute types of inconsistency and pairwise inconsistency, either inter-modality or intra-modality. Last but not least, we investigate some approaches to enhance the model performance in our probing experiment. The results indicate that text-based Chain-of-Thought prompting and visual-based prompting (Set-of-Mark annotations) offer minimal and sometimes adverse effects, whereas an iterative multimodal interleaved reasoning strategy shows promising gains. Overall, these results highlight a critical bottleneck in the ability of MLLMs to perform robust, integrated reasoning—a key challenge for future research.
@@
  • We perform a comprehensive evaluation of one leading multimodal reasoning model and five state-of-the-art MLLMs, revealing significant gaps in their ability to detect inconsistency errors with detailed error analyses across multiple error types, modalities, and layout complexities.
@@ 3 MMIR
A. Factual Contradiction: Direct conflict between two elements (text–text, text–image, or image–image) within the modified content.
@@
Figure 2 provides one example from each error type across web, office, and poster artifacts, illustrating the diverse challenges MMIR poses.
@@ 3.1 Data Curation
Finally, human experts review 747 edited samples, resulting in a final dataset of 534 validated quintuples: $D_{MMIR}=\{S^{\prime}_{i},E^{\prime}_{i},\mathtt{GT}_{i},\mathtt{category}_{i},\mathtt{rationale}_{i}\}_{i=1}^{534}$, ensuring that only realistic and challenging samples remain. Table 1 provides a detailed breakdown by artifact type, subcategory, and error type. For example, webpages are further divided into shopping, wiki, and classifieds, each with its average number of elements, while errors are distributed across the five defined categories. Notably, the average word count in multiple-choice questions is 382.6, whereas open-ended responses are fixed at 59 words.
@@ 4 Experiments and Analysis
We first evaluate the advanced multimodal reasoning model o1 OpenAI (2024b) and five other state-of-the-art MLLMs: GPT-4o OpenAI (2024a), Qwen2.5-VL Team (2025), LLaVA-NeXT Liu et al. (2024b), InternVL2.5 Chen et al. (2024) and Phi-3.5-Vision Abdin et al. (2024) on the MMIR benchmark. We implement open-source models using their default settings and select the 1217 version of o1 and the 1120 version of GPT-4o for evaluation. Model implementation details are provided in Appendix B. We then examine error patterns across different inconsistency types and layout complexities and finally explore how prompting strategies affect multimodal reasoning under the open-ended setting.
@@ 4.1 Main Results
As shown in Table 2, proprietary models (o1 and GPT-4o) significantly outperform open-source alternatives, though all models exhibit substantial room for improvement. Appendix A.4 shows a qualitative example with question-answer and model response.
Performance Gap Between Reasoning, Proprietary and Open-Source Models. In both open-ended and MCQ settings, the reasoning o1 model substantially outperforms the rest, surpassing all open-source models by over 30%. The other proprietary model GPT-4o, although missing the explicit reasoning ability of o1, outperforms open-source alternatives, reflecting stronger multimodal alignment and reasoning capabilities.
Impact of Semantic Cues. GPT-4o sees a 14.61% accuracy boost in the MCQ setting with additional element descriptions as options, narrowing its gap with o1 from 18.26% to just 4.4%. This indicates that GPT-4o relies heavily on semantic context when available.
Inconsistent Gains for Open-Source Models. Most open-source models gain moderate or little accuracy when provided with MCQ-style prompts. Phi-3.5-Vision-4B experiences a 9.93% drop, suggesting weaker reasoning capacity and less effective use of textual cues. The gap between proprietary and open-source models widens further in MCQ (from 27.08% to 35.21%), highlighting the persistent challenge of integrating perceptual grounding with logical inference.
@@ 4.2.1 Results Across Inconsistency Categories and Modalities
Figure 4(a) breaks down accuracy by the five inconsistency error categories. Proprietary models (o1, GPT-4o) outperform open-source models across the board, but the gap is particularly pronounced for Factual Contradictions and Identity Misattribution, implying that high-capacity models may have stronger factual grounding and entity recognition. Interestingly, Quantitative Discrepancy poses a substantial challenge for all models, highlighting a limitation in reasoning about numerical misalignment.
@@
In Figure 4(b), we examine how accuracy varies by the modality of the inconsistency. Overall, inter-modality errors yield the highest performance, with image-image inconsistencies proving especially tractable. Next in difficulty are intra-modality errors (image-text), which require partial cross-modal integration but can still leverage textual anchors. Finally, single modality (those involving only one text or image field) inconsistencies pose the greatest challenge, as they demand more advanced visual understanding and the ability to reconcile contextual inconsistency without contradiction with another element. These findings highlight that while models cope relatively well with pairwise conflicts, their capacity for deep visual or contextual reasoning remains underdeveloped.
@@ 4.3 Probing on Prompting Methods
We further investigate whether textual or visual prompts can alleviate the reasoning bottleneck. Table 3 compares Chain-of-Thought (CoT) prompting Wei et al. (2022) and Set-of-Mark (SoM) visual augmentation Yang et al. (2023), as well as their combination. We also explored an interleaved multimodal reasoning strategy, which we term Multimodal Interleaved CoT (MM-CoT) to further integrate and refine reasoning across both visual and textual modalities.
@@ 4.3.1 Chain-of-Thought (CoT) Prompting
To assess whether explicit reasoning instructions can enhance performance, we apply CoT prompting Wei et al. (2022) to the four open-sourced models (benchmarked proprietary models have API guides to not include additional CoT prompting).
@@ 4.3.2 Set-of-Mark (SoM) Prompting
We next examine the effect of SoM visual prompting Yang et al. (2023). By overlaying bounding boxes onto the artifact screenshots (example in Figure 6), we aim to enhance the models’ ability to perceive and localize elements.
@@ A.1 Inconsistency Error Category Definitions
    Direct conflict between two or more elements (text–text, text–image, or image–image) within the modified content.
@@ A.3 Auto-Verification and Editing Process
A.4 A Qualitative Example
@@
Appendix B Model Application Details
@@
For Qwen2.5-VL, we implemented the 7B version following the official repository: https://github.com/QwenLM/Qwen2.5-VL.
@@
Appendix C Data Release

### v3

@@ Abstract
Existing Multimodal Large Language Models (MLLMs) are predominantly trained and tested on consistent visual-textual inputs, leaving open the question of whether they can handle inconsistencies in real-world, layout-rich content. To bridge this gap, we propose the Multimodal Inconsistency Reasoning (MMIR) benchmark to assess MLLMs’ ability to detect and reason about semantic mismatches in artifacts such as webpages, presentation slides, and posters. MMIR comprises 534 challenging samples, each containing synthetically injected errors across five reasoning-heavy categories: Factual Contradiction, Identity Misattribution, Contextual Mismatch, Quantitative Discrepancy, and Temporal/Spatial Incoherence. We evaluate eight state-of-the-art MLLMs, showing that models with dedicated multimodal reasoning capabilities, such as o1, substantially outperform their counterparts while open-source models remain particularly vulnerable to inconsistency errors. Detailed error analyses further show that models excel in detecting pairwise inconsistencies but struggle with inconsistencies confined to single elements in complex layouts. Probing experiments reveal that single-modality prompting, including Chain-of-Thought (CoT) and Set-of-Mark (SoM) methods, yields marginal gains, revealing a key bottleneck in cross-modal reasoning. Our findings highlight the need for advanced multimodal reasoning and point to future research on multimodal inconsistency.
@@ 1 Introduction
Recent advances in Large Language Models (LLMs) have demonstrated impressive reasoning abilities across a variety of tasks OpenAI (2024b); Guo et al. (2025); Kojima et al. (2022); Wei et al. (2022). Building on pre-trained LLMs, Multimodal Large Language Models (MLLMs) are evolving rapidly. However, they usually face greater challenges as they need to reason across different modalities, especially when inconsistencies (i.e., mismatched or contradictory contents) exist. We find that, being primarily trained and evaluated on consistent visual-textual inputs, existing MLLMs are largely untested in scenarios where the input contains misaligned or contradictory information—a situation that is common in real-world scenarios. For example, in Figure 1, a user presents a web page containing conflicting visual and textual elements, asking the model to identify errors.
@@
To comprehensively evaluate the ability of MLLMs in reasoning over multimodal inconsistency, we introduce the Multimodal Inconsistency Reasoning Benchmark (MMIR). MMIR is the first framework dedicated to evaluating how effectively MLLMs can reason about and identify semantic mismatches within complex, layout-rich content with interleaved image and text components. Our benchmark is built on a diverse collection of real-world artifacts (e.g., websites, slides, posters) which have been augmented with synthetic inconsistencies—realistic inconsistency errors injected into their original structures. These inconsistency errors span a range of reasoning-heavy categories: Factual Contradiction, Identity Misattribution, Contextual Mismatch, Quantitative Discrepancy, and Temporal/Spatial Incoherence, posing a next-level reasoning challenge for models. For example, resolving a Identity Misattribution involves verifying entity alignment across modalities, while Quantitative Discrepancy requires cross-referencing chart data with textual claims. By challenging models to detect such inconsistencies, MMIR forces them to perform intricate reasoning that goes well beyond simple pattern recognition. Our benchmark not only exposes the limitations of current MLLMs in handling real-world challenges of reasoning over multimodal content with inconsistency, but also provides a platform for developing more robust multimodal reasoning systems.
@@
In our experiments, we evaluated the advanced multimodal reasoning model o1 OpenAI (2024b) and seven other state-of-the-art MLLMs: GPT-4o OpenAI (2024a), Qwen2.5-VL Team (2025), Llama-3.2 Grattafiori et al. (2024), LLaVA-NeXT Liu et al. (2024b), InternVL2.5 Chen et al. (2024), and Phi-3.5-Vision Abdin et al. (2024) using MMIR’s 534 test samples. The results overall underscore that current MLLM models struggle with multimodal inconsistency reasoning. Specifically, there is a stark contrast between proprietary and open-source models. The best open-source model evaluated only reaches less than 40% accuracy. o1 with strong reasoning capability achieves the overall best performance with over 50% accuracy.
@@
To further understand the benchmarking results, we conduct an analysis based on the inconsistency category, modality, and layout complexity of the artifact. We find the proprietary models excel in identifying factual contradiction and identity misattribute types of inconsistency and pairwise inconsistency, either inter-modality or intra-modality. Last but not least, we investigate some approaches to enhance the model’s performance in our probing experiment. The results indicate that text-based Chain-of-Thought prompting and visual-based prompting (Set-of-Mark annotations) offer minimal and sometimes adverse effects, whereas an iterative multimodal interleaved reasoning strategy shows promising gains. Overall, these results highlight a critical bottleneck in the ability of MLLMs to perform robust, integrated reasoning—a key challenge for future research.
@@
  • We perform a comprehensive evaluation of one leading multimodal reasoning model and seven state-of-the-art MLLMs, revealing significant gaps in their ability to detect inconsistency errors with detailed error analyses across multiple error types, modalities, and layout complexities.
@@ 3 MMIR
A. Factual Contradiction: Direct conflict between two elements (text–text, text–image, or image–image) within the artifact.
@@
Figure 2 provides one example from each error type across web, office, and poster artifacts, illustrating the diverse challenges MMIR poses. Detailed definitions and examples of inconsistency error types can be found in Appendix A.1.
@@ 3.1 Data Curation
Finally, human experts review 747 edited samples, resulting in a final dataset of 534 validated quintuples: $D_{MMIR}=\{A^{\prime}_{i},E^{\prime}_{i},\mathtt{GT}_{i},\mathtt{category}_{i},\mathtt{rationale}_{i}\}_{i=1}^{534}$, ensuring that only realistic and challenging samples remain. Table 1 provides a detailed breakdown by artifact type, subcategory, and error type. For example, webpages are further divided into shopping, wiki, and classifieds, each with its average number of elements, while errors are distributed across the five defined categories. Notably, the average word count in multiple-choice questions is 382.6, whereas open-ended responses are fixed at 59 words.
@@ 4 Experiments and Analysis
We first evaluate the advanced multimodal reasoning model o1 OpenAI (2024b) and seven other state-of-the-art MLLMs: GPT-4o OpenAI (2024a), Qwen2.5-VL Team (2025), Llama-3.2 Grattafiori et al. (2024), LLaVA-NeXT Liu et al. (2024b), InternVL2.5 Chen et al. (2024) and Phi-3.5-Vision Abdin et al. (2024) on the MMIR benchmark. We implement open-source models using their default settings and select the 1217 version of o1 and the 1120 version of GPT-4o for evaluation. Model implementation details are provided in Appendix C. We then examine error patterns across different inconsistency types and layout complexities, and finally explore how prompting strategies affect multimodal reasoning under the open-ended setting.
@@ 4.1 Main Results
As shown in Table 2, proprietary models (o1 and GPT-4o) significantly outperform open-source alternatives, though all models exhibit substantial room for improvement. Appendix B shows a qualitative example with question-answer and model response with qualitative analysis results.
Performance Gap Between Reasoning, Proprietary and Open-Source Models.
In both open-ended and MCQ settings, the reasoning o1 model substantially outperforms the rest, surpassing all 7B open-source models by over 24%. Qwen2.5-VL-72B performs best among the tested open-source models, with an overall accuracy of 27.24% in the vanilla setting. The other proprietary model, GPT-4o, although missing the explicit reasoning ability of o1, outperforms open-source alternatives, reflecting stronger multimodal alignment and reasoning capabilities.
Impact of Semantic Cues.
GPT-4o sees a 14.61% accuracy boost in the MCQ setting with additional element descriptions as options, narrowing its gap with o1 from 18.26% to just 4.4%. This indicates that GPT-4o relies heavily on semantic context when available. Similar accuracy boosts with additional semantic cues can be witnessed on Qwen2.5-VL-72B and Llama-3.2-90B-Vision-Instruct.
Inconsistent Gains for Open-Source Models.
Open-source models gain moderate or little accuracy when provided with MCQ-style prompts compared to proprietary GPT-4o. Phi-3.5-Vision-4B experiences a 9.93% drop, suggesting weaker reasoning capacity and less effective use of textual cues. The gap between proprietary and open-source models widens further in MCQ, highlighting the persistent challenge of integrating perceptual grounding with logical inference.
@@ 4.2.1 Results Across Inconsistency Categories and Modalities
Figure 4(a) breaks down accuracy by the five inconsistency error categories. Proprietary models (o1, GPT-4o) outperform open-source models across the board, but the gap is particularly pronounced for Factual Contradictions, implying that high-capacity models may have stronger factual grounding and entity recognition. Interestingly, Quantitative Discrepancy poses a substantial challenge for all models, highlighting a limitation in reasoning about numerical misalignment.
@@
In Figure 4(b), we examine how accuracy varies by the modality of the inconsistency. Overall, inter-modality errors yield the highest performance, with image-image inconsistencies proving especially tractable. Next in difficulty are errors involving textual elements, including intra-modality errors (image-text) and errors posed by inconsistent textual elements, which require partial cross-modal and cross-element integration but can still leverage textual anchors. Finally, single image modality (those involving only one image field) inconsistencies pose the greatest challenge, as they demand more advanced visual understanding and the ability to reconcile contextual inconsistency without contradiction with another element. These findings highlight that while models cope relatively well with pairwise conflicts, their capacity for deep visual or contextual reasoning remains underdeveloped.
@@ 4.3 Probing on Prompting Methods
We further investigate whether textual or visual prompts can alleviate the reasoning bottleneck. Table 3 compares Chain-of-Thought (CoT) prompting Wei et al. (2022) and Set-of-Mark (SoM) visual augmentation Yang et al. (2023), as well as their combination on six of the tested models. We also explored an interleaved multimodal reasoning strategy, which we term Multimodal Interleaved CoT (MM-CoT) to further integrate and refine reasoning across both visual and textual modalities.
@@ 4.3.1 Chain-of-Thought (CoT) Prompting
To assess whether explicit reasoning instructions can enhance performance, we apply CoT prompting Wei et al. (2022) to the four open-sourced models (benchmarked proprietary models have API guides to not include additional CoT prompting). CoT prompting is a technique that encourages models to solve complex problems by generating intermediate reasoning steps, thereby enhancing their problem-solving abilities. Inspired by the common CoT setup, we append a text prompt: “Think step by step first, then provide the final answer.” to each base input in this ablation setting.
@@ 4.3.2 Set-of-Mark (SoM) Prompting
We next examine the effect of SoM visual prompting Yang et al. (2023). SoM prompting is a visual prompting method that enhances the visual grounding abilities of LMMs by overlaying images with spatial and speakable marks, such as alphanumerics or boxes, to partition the image into regions at different levels of granularity. Following the original work, we use visual annotations (e.g., bounding boxes with numerical IDs) to highlight elements in the artifact, as shown in Figure 6. This is a visual analog to CoT, aiming to guide the model’s attention during inference.
@@ A.1 Inconsistency Error Category Definitions
    Direct conflict between two or more elements (text–text, text–image, or image–image) within the artifact.
@@ A.3 Auto-Verification and Editing Process
Appendix B Qualitative Example and Analysis
Figure 7 illustrates a test sample with model responses under the two main settings in MMIR: open-ended and multiple-choice.
Beyond the single illustration, we performed a qualitative analysis of the top-performing models’ (o1, GPT-4o) performance in the open-ended format, examining an additional 50 randomly selected samples. Our analysis revealed multiple common failure modes:
  • Contextual misunderstanding: The model frequently flagged false positives when required to integrate contextual information, such as distinguishing between user-selected and available options. For instance, on a shopping website listing a “Bath Body Brush,” with a photo showing a pink brush and selectable color options (Pink, Green, etc.), the model incorrectly flagged inconsistency due to confusion over color selection context (the photo and the unselected option of “Green”), despite clear visual alignment between the image and selected color option “Pink”.
  • Indirect semantic reasoning: Errors often arose when inconsistencies required inference beyond direct semantic contradictions. The model struggled notably with subtle contradictions, such as mismatches between implied and explicit information (e.g., temporal contexts like "limited-time offers" vs. static pricing information).
  • Complex textual layouts: The model’s accuracy significantly decreased when inconsistencies were embedded within lengthy or dense textual sections, such as detailed paragraphs in product descriptions, compared to simpler fields like product titles.
@@ Appendix B Qualitative Example and Analysis
Appendix C Model Application Details
@@
For Qwen2.5-VL, we implemented the 72B and 7B versions following the official repository: https://github.com/QwenLM/Qwen2.5-VL.
For Llama-3.2, we utilized the API following the API guidelines available at https://openrouter.ai/qwen/qwen-2.5-72b-instruct.
@@
Appendix D Data Release